5(2-x)/6 = x+6/2 Trying to teach myself algebra from text. If you would do this one step by step, I would really appreciate it. Thanks
Answer by stanbon(75887)

Since you have denominators, find the LCM (least common multiple) of 6 and 2.
The LCM is 6.
Multiply both sides of the equation by the LCM = 6 to get the following:
5(2-x) = 6x +18
10-5x = 6x+18
Add 5x to both sides and subtract 18 from both sides to get:
11x = -8
Divide both sides by 11 to get
x = -8/11